OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (bradely):

Slope of the line is -1 Perpendicular line slope ,m =-1/-1=1 y=x+c Point (-4,-2) -2=-4+c c=2 y=x+2 option 4 is correct http://www.mathskey.com/question2answer/ask

OpenStudy (anonymous):

Thank You bradley, can you help me with one more.